# Frustratingly Easy Zero-Day Audio DeepFake Detection via Retrieval Augmentation and Profile Matching

### v1

@@ 4.2 Ablation study
The effect of voice profile informative attributes. We use S08 to create and compare variants by probing the attributes in the profile features: age & gender, emotion, and voice quality. We remove each attribute individually to assess its contribution to performance. Results are presented in Table 2. While properties more closely related to speakers, such as age, sex and emotion, have a relatively negligible impact, removing voice quality substantially deteriorates the performance, especially for shorter segments. Considering the nature of the DE2024, this indicates the relatively higher importance of voice qualities for the retrieval, which can be affected by not only the speaker, but also the surrounding environment and context.
Fig. 3: The number of retrieved utterances and the performance of the proposed in EER. The red and green dashed lines denotes the baseline and fine-tuned system performance in Table 1, respectively. Best viewed in color.
4s 30s
@@
S08 14.98 82.63 11.42 87.13
w/o Age & Gender 14.86 82.67 13.86 85.65
w/o Emotion 14.86 82.47 12.57 86.31
w/o Voice Quality 27.20 78.70 13.71 83.15
Table 2: Probing attributes from profile feature extractor (w/o: without). Number of retrieved utterances here is $k=10$.
5 Conclusion
@@
In this study, we addressed zero-day attacks in audio deepfake detection, where systems must use existing knowledge to rapidly handle unseen attacks. We have introduced an effective strategy based on knowledge databases and retrieval augmentation, with various retrieval and ensemble methods, without any manner of fine-tuning or additional training. The framework uses both state-of-the-art CM and voice profile representations. Leveraging DE2024 with various length of segments, we found that the proposed method, especially when using both representations and statistics of retrieved data directly for ensemble, achieves competitive performance against fine-tuned CM models. Future work may explore acquiring temporal information via frame-level SSL features, leveraging profile information more effectively, and extending the framework to multi-classification ADD-related tasks.

### v2

@@ IV-B Ablation on voice profile attributes
Results are presented in Table II. While properties more closely related to speakers, such as age, sex and emotion, have a relatively weaker impact, removing voice quality more substantially deteriorates the performance in both metrics, especially for 4-second condition. This indicates the relatively higher importance of voice qualities, which can be affected by not only the speaker, but also the surrounding environment and context, considering the nature of the DE2024.
V Cross-database evaluation
In the previous sections, the knowledge retrieval source, fine-tuning data, and evaluation sets were all derived from DE2024. Despite its in-the-wild nature, this “in-domain” setting does not fully show how well the proposed RA generalize to other databases.
Therefore, in this section, we keep the knowledge source for RA (and fine-tuning data) fixed to DE2024 (4-second condition), and additionally include AI4T as part of evaluation, as mentioned in Section III-B. We further extend the RA framework by testing two simple fusion methods that may improve cross-database generalizability. We also evaluate several previously proposed countermeasures [9, 17], together with their fine-tuned variants, where fine-tuning is performed on the DE2024 training data.
V-A Linear and Selective fusion
From the results shown in Table I, it was found that the proposed method (e.g. S08) is training-free and holds promising performance, while the fine-tuned system still performs moderately better. These two approaches may provide complementary strengths.
Therefore, we evaluate two simple fusion strategies here: 1) Linear fusion, where we consider simple fusion between $s_{q}$ and $\hat{s}_{q}$ shown in Section II-C by $s_{linear}=\lambda*s_{q}+(1-\lambda)*\hat{s}_{q}$; 2) Selective fusion, where the query representation is firstly input to a specialized module for deciding whether it is in-domain or out of domain (OOD), then routed to the most suitable detector. If the query is judged as in-domain, we directly use the score from the SSL-based CM. Conversely, if the query is judged to be OOD, we switch to the proposed RA method with its knowledge database, and use the ensembled score as the final output.
We employ a training-free $k$-NN based OOD that measures cosine similarity between input embedding and one in a knowledge database [14]. For better understanding, the framework of selective fusion is illustrated in Fig. 3. Here, we utilize two knowledge databases with distinct roles: 1) A knowledge database dedicated to the training-free OOD detection; 2) Another database used for RA of the proposed deepfake detector.
Figure 3: Selective fusion of the proposed training-free deepfake detector and SSL-based CM. The lock symbols indicate the models that are remain frozen without fine-tuning. The text in italian refers to the source data to construct the knowledge databases. We use B00 and S08 as example systems here.
DE2024 (4 seconds) AI4T
@@ V-A Linear and Selective fusion
No fine-tuning with DE2024 training data
AASIST [9] - 55.73 37.71 - 55.81 57.30
P3 [17] - 47.83 56.15 - 53.45 56.26
B00 - 27.76 75.15 - 10.48 75.51
S08 - 14.78 86.84 - 16.98 83.24
B00 ++ S08 - 13.78 87.20 - 15.92 84.06
B00 $\lor$ S08 0.76 14.22 87.16 1.00 16.98 83.24
Fine-tuned with with DE2024 training data
AASIST - 18.87 82.50 - 53.91 44.56
P3 - 14.04 86.07 - 17.91 67.92
B01 - 12.52 87.69 - 8.94 90.20
S08* - 12.56 89.11 - 14.59 85.64
B01 ++ S08* - 12.02 89.65 - 13.29 87.12
B01 $\lor$ S08* 0.62 12.25 89.19 0.43 11.71 89.02
TABLE III: Results on the cross-database evaluation. $\mathcal{R}$ denotes the proportion of OOD data classified from the query dataset (1.00 means all data are OOD). Numbers in bold indicate the best results.
V-B Implementation details
Data. Since they include both outputs from SSL-based CM and RA, we consider two system pairs: 1) B00 and S08, which are based on the SSL-based CM without fine-tuning; 2) B01 and S08*, which are from the SSL CM fine-tuned on DE2024 training data (S08* represents RA built on B01). In terms of the source data of the two knowledge databases, the knowledge database for the OOD detection is constructed from a gender- and language-balanced subsample of the original training data used to post-train the SSL-based CM [^7] Details of the training data, including statistics and attribute-wise distributions, can be found in https://github.com/nii-yamagishilab/AntiDeepfake.. The size of the database is 26,000 samples, which is almost same as the DE2024 4-second training data. To set the OOD decision threshold, we further construct a small development set by randomly sampling 1,000 utterances from either knowledge database. Among the RA configurations, we use S08 with the knowledge database described in Section II-C.
Implementation. For linear fusion, we need to determine the weight value $\lambda$ between the SSL CM scores and the RA ensembled scores on DE2024 development set, and use such $\lambda$ for all query datasets involved. In this section, $\lambda=0.1$ for both fusion methods. For the $k$-NN based OOD detection, $k$ was set to 20. The linear and selective fusion methods of two systems $A$ and $B$ are correspondingly denoted with symbols $A+B$ and $A\lor B$ in Table III, respectively.
V-C Results
The results are summarized in Table III. We analyze the results separately depending on whether the backbone SSL-based CM model is fine-tuned with the DE2024 training data or not.
No fine-tuning with DE2024 training data. On the DE2024 evaluation data, S08 substantially improves over the baseline B00 and outperforms prior works [9, 17]. The linear fusion B00++S08 resulted in the best EER and accuracy. However, on the more-mismatched AI4T, S08 exhibits rather not promising zero-shot generalization in terms of EER compared to the direct application of the SSL-based CM. This indicates that the proposed RA-based deepfake detection method (S08) is training-free, but suffers from the domain mismatch between the knowledge database based on the DE2024 training data and AI4T. The linear fusion improves the accuracy over using S08 alone, but yields worse EER.
Fine-tuning with DE2024 training data. Even if the models are fine-tuned with the DE2024 training data, the general tendency is the same as the above non fine-tuning case. On the DE2024 evaluation, S08 achieves performance comparable to the fine-tuned SSL-based CM B01 and outperforms prior works [9, 17]. The linear fusion B01 ++ S08* achieves the best overall performance on DE2024, reaching the lowest EER (12.02%) and highest accuracy (89.65%). Meanwhile, on AI4T, S08 is not the good choice anymore and the baseline B01 resulted in the best performance in terms of both the accuracy and EER. These observations indicate that even training-free approaches leveraging knowledge databases may be affected by domain mismatch, and addressing this will likely be an interesting and important research direction for future work.
VI Conclusion
@@
In this study, we have addressed zero-day attacks in audio deepfake detection, where systems must use existing knowledge to rapidly handle unseen attacks. We have introduced an retrieval augmentation framework based on knowledge databases and various related strategies, without any manner of fine-tuning or additional training. The framework uses both state-of-the-art CM and voice profile representations. Leveraging DE2024 with various duration conditions, we have found that the proposed method, especially when combining SSL and profile features from the retrieved data, achieves competitive performance. Finally, we have evaluated the framework on the cross-database scenario and show that the even training-free approach leveraging knowledge databases may be affected by domain mismatch. Future work may explore acquiring temporal information via frame-level SSL features, modeling profile information more effectively, and extending the framework to multi-classification ADD-related tasks.